OpenStudy (anonymous):

Let f(x)=x^2+4. Find f(x+h)−f(x) h if h≠0 . Simplify your answer.

OpenStudy (zzr0ck3r):

\(\frac{(x+h)^2+4-x^2-4}{h}=\frac{x^2+2xh+h^2+4-x^2-4}{h}=\frac{2xh+h^2}{h}=2x+h\)